$SPAD/src/input de2re.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 bugs

The function —LinearOrdinaryDifferentialOperator— is undefined — bug1 —

dn := D()$L
-- 4 well known differential operators
exp:= dn - 1
sincos := dn**2 + 1
airy := dn**2 - n
hermite := dn**2 - 2*n*dn + 1
-- the recurrences satisfied by the coefficients of their series solutions
--recurrence(exp, 0)
--recurrence(sincos, 0)
--recurrence(airy, 0)
--recurrence(airy, 1)
--recurrence(hermite, 0)
-- a non-trivial example from the GFUN tech. rep (Salvy & Zimmermann)
op := (335 * n**2 + 1290) * dn**2 + 1540 * n * dn + 468720
--recurrence(op, 0)

———-

— bugs —

---------------------------------- de2re.input -----------------------------
)cl all
P := UP(n, INT)
n := n::P
L := LODO1 P
\getchunk{bug1}

———-

2 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
---------------------------------- de2re.input -----------------------------
)cl all
P := UP(n, INT)
n := n::P
L := LODO1 P

———-

References^{[1] nothing}

5